# Dancing in Chains: Strategic Persuasion in Academic Rebuttal via Theory of Mind

### v1

@@ Abstract
Although artificial intelligence (AI) has become deeply integrated into various stages of the research workflow and achieved remarkable advancements, academic rebuttal remains a significant and underexplored challenge. This is because rebuttal is a complex process of strategic communication under severe information asymmetry rather than a simple technical debate. Consequently, current approaches struggle as they largely imitate surface-level linguistics, missing the essential element of perspective-taking required for effective persuasion. In this paper, we introduce RebuttalAgent, the first framework to ground academic rebuttal in Theory of Mind (ToM), operationalized through a ToM-Strategy-Response (TSR) pipeline that models reviewer mental state, formulates persuasion strategy, and generates strategy-grounded response. To train our agent, we construct RebuttalBench, a large-scale dataset synthesized via a novel critique-and-refine approach. Our training process consists of two stages, beginning with a supervised fine-tuning phase to equip the agent with ToM-based analysis and strategic planning capabilities, followed by a reinforcement learning phase leveraging the self-reward mechanism for scalable self-improvement. For reliable and efficient automated evaluation, we further develop Rebuttal-RM, a specialized evaluator trained on over 100K samples of multi-source rebuttal data, which achieves scoring consistency with human preferences surpassing powerful judge GPT-4.1. Extensive experiments show RebuttalAgent significantly outperforms the base model by an average of 18.3% on automated metrics, while also outperforming advanced proprietary models across both automated and human evaluations. Disclaimer: the generated rebuttal content is for reference only to inspire authors and assist in drafting. It is not intended to replace the author’s own critical analysis and response. [^1] Our code and models: https://github.com/Zhitao-He/RebuttalAgent
@@ 1 Introduction
In this paper, we propose RebuttalAgent, the first model to integrate Theory of Mind into academic rebuttal. RebuttalAgent employs a novel three-stage generation framework we term ToM-Strategy-Response (TSR), which decomposes the complex task of rebuttal into a coherent series of reasoning and generation steps. Specifically, the initial Theory-of-Mind (T) stage comprises a hierarchical analysis to discern macro-level reviewer intent while deconstructing the micro-level attributes of each comment. This analysis constructs a multi-dimensional reviewer profile designed to inform both global strategy and local tactics. Subsequently, the Strategy (S) stage utilizes this profile to formulate an actionable plan for the target comment, which aligns the response strategy with both the macro- and micro-level critiques from the reviewer. The process concludes with the Response (R) stage, which achieves context-aware synthesis by integrating the reviewer profile, the plan, and pre-retrieved evidential chunks from the original manuscript, thereby generating a persuasive response.
@@
  • We introduce RebuttalAgent, the first framework to leverage Theory of Mind (ToM) for academic rebuttal. Our agent employs a novel ToM-Strategy-Response (TSR) pipeline. By explicitly modeling the reviewer’s perspective, identifying key concerns, and suggesting grounded responses with adaptive strategic reasoning, our agent aims to help authors communicate more clearly and effectively and move beyond formulaic responses.
  • We construct RebuttalBench, a large-scale dataset of over 70K high-quality samples created via our critique-and-refine pipeline, with each sample containing a ToM-strategy-response chain. Building on the foundational ToM-based reasoning and rebuttal capabilities through SFT, we further optimize the analysis and strategic policies of agent using RL with our Self-reward mechanism, enabling scalable policy refinement without external reward model.
  • To conduct reliable and efficient evaluation, we develop Rebuttal-RM, a specialized evaluator that achieved high scoring consistency with human experts. Extensive experiments show RebuttalAgent outperforms the base model by an average of 18.3%, and shows performance comparable to advanced proprietary across both automated and human evaluation.
@@ 2 Task Formulation
Figure 1: Overview of our RebuttalAgent framework. First, we extract each comment from raw reviews and retrieves their relevant context from the paper. Next, based on our TSR pipeline, we collect a tailored strategy and response for each comment, grounded in Theory of Mind. Finally, our RebuttalAgent is trained via Supervised Fine-Tuning, followed by Reinforcement Learning with a self-reward mechanism, enabling both scalability and self-improvement.
@@ 3.2 Context Retrieval
A single reviewer comment typically targets a specific aspect of the manuscript, such as a formula or baseline comparison. However, using the full, information-dense manuscript as context is infeasible and sub-optimal, as it can overwhelm the model and dilute focus. Therefore, we implement a three-stage context retrieval pipeline to isolate the most relevant content for each comment. As shown in the top-left corner of Figure 1, the retrieval pipeline begins by segmenting the manuscript ($M$) into discrete text chunks, typically corresponding to paragraphs. Then we employ a pre-trained embedding model [^2] https://huggingface.co/Qwen/Qwen3-Embedding-0.6B to encode both the target comment ($c_{\text{target}}$) and each text chunk into high-dimensional vector representations. Relevance is then quantified by computing the cosine similarity between the comment vector and all chunk vectors. Finally, the top-$k$ chunks with the highest similarity scores are retrieved to serve as the context. The effectiveness analysis of retrieval module is provided in Appendix B.
@@ 4 ToM-Strategy-Response Framework
Theory of Mind (ToM) is a core concept in cognitive science, referring to the ability to understand and reason about the differing beliefs, intentions, desires, and perspectives of others. Applying this concept to artificial intelligence has led to Machine Theory of Mind (MToM), which is an AI system’s capacity to infer and model the mental states of human or AI teammates to support collaboration. Large language models such as GPT-4 have demonstrated stronger ToM-like reasoning capabilities. In our work, we extend MToM to the specific domain of academic rebuttal. Given the game-theoretic and information-asymmetric nature of the rebuttal process, modeling the reviewer’s beliefs, knowledge background, and core concerns is particularly critical. Therefore, our proposed RebuttalAgent framework explicitly implements ToM through a Theory-of-Mind-Strategy-Response (TSR) pipeline, which first constructs a hierarchical reviewer profile to guide the subsequent formulation of strategy and response. Figure 1 (bottom) depicts how our RebuttalAgent framework decomposes the task of rebuttal into a multi-stage reasoning process: (1) inferring the reviewer’s perspective with ToM, (2) formulating a tailored strategy, and (3) synthesizing a convincing, evidence-grounded response.
@@ 4.2 ToM-Driven Strategy Generation
4.3 Strategy-Guided Refined Response Generation
The final stage of our TSR pipeline generates the definitive response ($r_{\text{target}}$) through an advanced guided synthesis process, conditioned on a rich set of strategic and contextual inputs. This intricate process is informed by two distinct yet complementary primary types of input:
@@ 7.1 Evaluation of Rebuttal-RM
To validate the effectiveness of Rebuttal-RM, we conduct comprehensive evaluation to measure the agreement between our model and human experts. Following recent work (Wu et al., 2025), we employ a set of six statistical metrics. We use four standard statistical measures to assess the overall correlation: Mean Absolute Error ($e$), Pearson ($r$), Spearman ($\beta$), and Kendall ($\tau$). Additionally, to mitigate potential annotator biases and assess classification accuracy, we introduce two metrics based on score ranges: Coarse-grained Accuracy ($c$): Scores are mapped to four quality tiers: Unconvincing (scores 1-3), Acceptable (scores 4-6), Good (scores 7-8), and Excellent (scores 9-10). Fine-grained Accuracy ($f$): For a stricter assessment, scores are categorized into seven more granular ranges derived from our rubric, such as grouping scores of 1 and 2, 3 and 4, and so on, with single-point ranges for scores of 5 and 6.
@@ 7.3 Experimental Results
Ablation Study. Our ablation study confirms the necessity of all the model’s design components. Performance significantly drops when removing any key component, such as ToM, Strategy, Thinking, or when omitting core training stages such as SFT and RL. Among all reward signals, the one for final response quality proved to be the most impactful. These results show that our model’s success is rooted in the synergy between its specialized data, complete training process, and reward mechanism. Applying our framework to Llama-3.1-8B and Qwen3-4B yields significant gains, raising scores from 7.44 to 9.20 and 7.69 to 8.98, respectively. These results demonstrate that the effectiveness of our TSR pipeline and self-reward mechanism is not tied to a specific backbone; rather, it serves as a model-agnostic strategy that generalizes well to other models, including smaller ones.
@@ 8 Related Work
Machine Theory of Mind. Machine Theory of Mind (ToM) refers to an AI system’s capacity to infer and model the mental states of human or AI teammates to support collaboration (Rabinowitz et al., 2018; Goldman and others, 2012; Wellman, 2002; Yang et al., 2025c; Leslie et al., 2004). Instruction-tuned models such as GPT-4 have demonstrated stronger ToM-like reasoning compared to earlier versions (Kosinski, 2023; 2024), sometimes matching or exceeding human performance in tasks involving sarcasm and social inference. Various methods have been proposed to explicitly model ToM. For example, SymbolicToM builds symbolic belief graphs to track character beliefs for answer generation (Sclar et al., 2023). SimToM employs perspective-taking and context filtering in a two-stage process (Wilf et al., 2023), while ToM-LM translates questions into symbolic forms for model checking (Tang and Belle, 2024). ToMAP integrates opponent modeling and reinforcement learning to generate more persuasive arguments (Han et al., 2025).
@@
LLM Debate. The use of multi-agent debate and interaction among Large Language Models (LLMs) has emerged as a promising approach to enhance capabilities in complex reasoning (He et al., 2023; 2024b; Qin et al., 2025; Xu et al., 2024; Yang et al., 2025b; Chen et al., 2025c) and fact-checking by simulating collaborative or adversarial dialogue (Du et al., 2023; He et al., 2025a; Liang et al., 2023; Jin et al., 2024a; Breum et al., 2024; He et al., 2025c; Salvi et al., 2025). For instance, ChatEval employs a multi-agent referee team to evaluate open-ended responses (Chan et al., 2023), while AgentsCourt improves answer quality through multi-round debate among model instances (He et al., 2024a). Debatrix provides a structured judging framework to assess debates along multiple dimensions (Liang et al., 2024), and DyLAN dynamically assembles agent teams tailored to different tasks (Liu et al., 2024c). Notably, Salvi et al. (2025) shows that GPT-4 equipped with sociodemographic data can outperform humans in persuasion.
@@
LLM for Academic Peer Review. The emerging paradigm of AI for Research applies Large Language Models (LLMs) to automate and enhance scholarly activities, including automated research (Schmidgall and Moor, 2025; Yamada et al., 2025) and writing assistance (Wang et al., 2025; Chen et al., 2025a). Within the critical domain of peer review, LLMs are leveraged for generating reviews (Zhu et al., 2025; Idahl and Ahmadi, 2025) and for enhancing review quality analysis (Purkayastha et al., 2025). Furthermore, multi-agent systems have been proposed to explore peer review dynamics (Jin et al., 2024b; D’Arcy et al., 2024) and automate research workflows (Schmidgall et al., 2025). Despite the creation of large, multi-turn review datasets (Zhang et al., 2025), there remains limited exploration into the rebuttal stage. Building on these foundations, our work proposes a RebuttalAgent framework that explicitly leverages Theory of Mind to model reviewer intent, enabling more strategic and context-aware responses.
@@ 9 Conclusion
In this paper, we introduce RebuttalAgent, the first framework to ground academic rebuttal in Theory of Mind (ToM). To train our agent, we construct RebuttalBench, a large-scale synthetic dataset created via a novel critique-and-refine pipeline. Our twofold training process begins with a Supervised Fine-tuning phase to equip the agent with ToM-based analysis and strategic planning capabilities, followed by a Reinforcement Learning phase using a novel self-reward mechanism. For a reliable and scalable automated evaluation, we develop Rebuttal-RM, a specialized evaluator trained on over 100K samples of multi-source data. Extensive experiments show RebuttalAgent significantly outperforms the base model by 18.3% and is competitive with advanced models such as o3 across both automated and human evaluations.
@@
This paper introduces a comprehensive framework for leveraging Theory of Mind (ToM) for academic rebuttal. This framework comprises three main components: (1) a rebuttal evaluator, Rebuttal-RM; (2) a large-scale high-quality dataset, RebuttalBench; and (3) a novel academic assistant, RebuttalAgent. To ensure the full reproducibility of this framework, we have provided detailed documentation across the paper and its appendices. The generation process for the RebuttalBench dataset, along with the complete training procedures for RebuttalAgent (including all hyperparameters), are provided in Section 5. The details for training Rebuttal-RM, the generation process for RAR-Rebuttal dataset are provided in Section 6. Our code and models will be released publicly for future research.

### v3

@@ Abstract
Although artificial intelligence (AI) has become deeply integrated into various stages of the research workflow and achieved remarkable advancements, academic rebuttal remains a significant and underexplored challenge. This is because rebuttal is a complex process of strategic communication under severe information asymmetry rather than a simple technical debate. Consequently, current approaches struggle as they largely imitate surface-level linguistics, missing the essential element of perspective-taking required for effective persuasion. In this paper, we introduce RebuttalAgent, the first framework to ground academic rebuttal in Theory of Mind (ToM), operationalized through a ToM-Strategy-Response (TSR) framework that models reviewer mental state, formulates persuasion strategy, and generates evidence-based response. To train our agent, we construct RebuttalBench, a large-scale dataset synthesized via a novel critique-and-refine approach. Our training process consists of two stages, beginning with a supervised fine-tuning phase to equip the agent with ToM-based analysis and strategic planning capabilities, followed by a reinforcement learning phase leveraging the self-reward mechanism for scalable self-improvement. For reliable and efficient automated evaluation, we further develop Rebuttal-RM, a specialized evaluator trained on over 100K samples of multi-source rebuttal data, which achieves scoring consistency with human preferences surpassing powerful judge GPT-4.1. Extensive experiments show RebuttalAgent significantly outperforms the base model by an average of 18.3% on automated metrics, while also outperforming advanced proprietary models across both automated and human evaluations. Disclaimer: the generated rebuttal content is for reference only to inspire authors and assist in drafting. It is not intended to replace the author’s own critical analysis and response. [^1] Our code and models: https://github.com/Zhitao-He/RebuttalAgent
@@ 1 Introduction
In this paper, we propose RebuttalAgent, the first model to integrate Theory of Mind into academic rebuttal. RebuttalAgent employs a novel three-stage generation framework we term ToM-Strategy-Response (TSR), which decomposes the complex task of rebuttal into a coherent series of reasoning and generation steps. Specifically, the initial Theory-of-Mind (T) stage comprises a hierarchical analysis to discern macro-level reviewer intent while deconstructing the micro-level attributes of each comment. This analysis constructs a multi-dimensional reviewer profile designed to inform both global strategy and local tactics. Subsequently, the Strategy (S) stage utilizes this profile to formulate an actionable plan for the target comment, which aligns the response strategy with both the macro- and micro-level critiques from the reviewer. The process concludes with the Response (R) stage, which achieves evidence-based synthesis by grounding the strategic plan in pre-retrieved evidential chunks to address the inferred reviewer concerns, thereby generating a persuasive response.
@@
  • We introduce RebuttalAgent, the first model to leverage Theory of Mind (ToM) for academic rebuttal, transforming the process from simple linguistic imitation into a strategic reasoning task. Our agent employs a novel ToM-Strategy-Response (TSR) framework that explicitly models the reviewer’s perspective, identifies underlying concerns, and formulates evidence-based responses through adaptive strategic reasoning. By bridging the gap between perception and persuasion, RebuttalAgent enables academic communication more clear, constructive.
  • We construct RebuttalBench, a large-scale dataset of over 70K high-quality samples created via our critique-and-refine pipeline, with each sample containing a ToM-strategy-response chain. Building on the foundational ToM-based reasoning and rebuttal capabilities through SFT, we further optimize the analysis and strategic policies of the agent using RL with our Self-reward mechanism, which allowing the agent to continuously improve its persuasive depth and strategic alignment without the need for an external, expert-annotated reward model.
  • We develop Rebuttal-RM, a specialized evaluator trained on over 100K multi-source samples, achieving a high degree of alignment with human expert preferences, significantly surpassing the consistency of powerful proprietary models like GPT-4.1. Extensive experiments show RebuttalAgent outperforms the base model by an average of 18.3%, and demonstrates performance comparable to advanced proprietary models across both automated and human evaluations.
@@ 2 Task Formulation
Figure 1: Overview of our RebuttalAgent framework. First, we extract each comment from raw reviews and retrieves their relevant context from the paper. Next, based on our TSR framework, we collect a tailored strategy and response for each comment, grounded in Theory of Mind. Finally, our RebuttalAgent is trained via Supervised Fine-Tuning, followed by Reinforcement Learning with a self-reward mechanism, enabling both scalability and self-improvement.
@@ 3.2 Context Retrieval
A single reviewer comment typically targets a specific aspect of the manuscript, such as a formula or baseline comparison. However, using the full, information-dense manuscript as context is infeasible and sub-optimal, as it can overwhelm the model and dilute focus. Therefore, we implement a three-stage context retrieval module to isolate the most relevant content for each comment. As shown in the top-left corner of Figure 1, the retrieval pipeline begins by segmenting the manuscript ($M$) into discrete text chunks, typically corresponding to paragraphs. Then we employ a pre-trained embedding model [^2] https://huggingface.co/Qwen/Qwen3-Embedding-0.6B to encode both the target comment ($c_{\text{target}}$) and each text chunk into high-dimensional vector representations. Relevance is then quantified by computing the cosine similarity between the comment vector and all chunk vectors. Finally, the top-$k$ chunks with the highest similarity scores are retrieved to serve as the context. The effectiveness analysis of retrieval module is provided in Appendix B.
@@ 4 ToM-Strategy-Response Framework
Theory of Mind (ToM) is a core concept in cognitive science (Leslie et al., 2004; Goldman and others, 2012), referring to the ability to understand and reason about the differing beliefs, intentions, desires, and perspectives of others. Applying this concept to artificial intelligence has led to Machine Theory of Mind (MToM) (Rabinowitz et al., 2018), which is an AI system’s capacity to infer and model the mental states of human or AI teammates to support collaboration. Large language models such as GPT-4 have demonstrated stronger ToM-like reasoning capabilities. In our work, we extend MToM to the specific domain of academic rebuttal. Given the game-theoretic and information-asymmetric nature of the rebuttal process, modeling the reviewer’s beliefs, knowledge background, and core concerns is particularly critical. Therefore, our proposed RebuttalAgent framework explicitly implements ToM through a Theory-of-Mind-Strategy-Response (TSR) framework, which first constructs a hierarchical reviewer profile to guide the subsequent formulation of strategy and response. Figure 1 (bottom) depicts how our RebuttalAgent framework decomposes the task of rebuttal into a multi-stage reasoning process: (1) inferring the reviewer’s perspective with ToM, (2) formulating a tailored strategy, and (3) synthesizing a convincing, evidence-grounded response.
@@ 4.2 ToM-Driven Strategy Generation
4.3 Evidence-Based Strategic Response Generation
The final stage of our TSR framework generates the definitive response ($r_{\text{target}}$) through an advanced guided synthesis process, conditioned on a rich set of strategic and contextual inputs. This intricate process is informed by two distinct yet complementary primary types of input:
@@ 7.1 Evaluation of Rebuttal-RM
Following recent work (Wu et al., 2025), we employ a set of six statistical metrics to validate the effectiveness of Rebuttal-RM. We use four standard statistical measures to assess the overall correlation: Mean Absolute Error ($e$), Pearson ($r$), Spearman ($\beta$), and Kendall ($\tau$). Additionally, to mitigate potential annotator biases and assess classification accuracy, we introduce two metrics based on score ranges: Coarse-grained Accuracy ($c$): Scores are mapped to four quality tiers: Unconvincing (scores 1-3), Acceptable (scores 4-6), Good (scores 7-8), and Excellent (scores 9-10). Fine-grained Accuracy ($f$): For a stricter assessment, scores are categorized into seven more granular ranges derived from our rubric, such as grouping scores of 1 and 2, 3 and 4, and so on.
@@ 7.3 Experimental Results
Ablation Study. Our ablation study confirms the necessity of all the model’s design components. Performance significantly drops when removing any key component, such as ToM, Strategy, Thinking, or when omitting core training stages such as SFT and RL. Among all reward signals, the one for final response quality proved to be the most impactful. These results show that our model’s success is rooted in the synergy between its specialized data, complete training process, and reward mechanism. Applying our framework to Llama-3.1-8B and Qwen3-4B yields significant gains, raising scores from 7.44 to 9.20 and 7.69 to 8.98, respectively. These results demonstrate that the effectiveness of our TSR framework and self-reward mechanism is not tied to a specific backbone, it serves as a model-agnostic strategy that generalizes well to other models, including smaller ones.
@@ 8 Related Work
Machine Theory of Mind. Machine Theory of Mind (ToM) refers to an AI system’s capacity to infer and model the mental states of human or AI teammates to support collaboration (Rabinowitz et al., 2018; Goldman and others, 2012; Wellman, 2002; Huang et al., 2025; Yang et al., 2025c; Leslie et al., 2004). Instruction-tuned models such as GPT-4 have demonstrated stronger ToM-like reasoning compared to earlier versions (Kosinski, 2023; 2024), sometimes matching or exceeding human performance in tasks involving sarcasm and social inference. Various methods have been proposed to explicitly model ToM. For example, SymbolicToM builds symbolic belief graphs to track character beliefs for answer generation (Sclar et al., 2023). SimToM employs perspective-taking and context filtering in a two-stage process (Wilf et al., 2023), while ToM-LM translates questions into symbolic forms for model checking (Tang and Belle, 2024). ToMAP integrates opponent modeling and reinforcement learning to generate more persuasive arguments (Han et al., 2025). In this work, we extend machine ToM to the academic rebuttal and propose RebuttalAgent, which operationalizes ToM through a hierarchical analysis of reviewer intent and a multi-stage reasoning pipeline.
@@
LLM Debate. The use of multi-agent debate and interaction among Large Language Models (LLMs) has emerged as a promising approach to enhance capabilities in complex reasoning (He et al., 2023; Xu et al., 2024b; He et al., 2024a; Ju et al., 2026; Yan et al., 2025; Shen et al., 2025; Qin et al., 2025; Xu et al., 2024a; Yang et al., 2025b; He et al., 2025d; Chen et al., 2025c) and fact-checking by simulating collaborative or adversarial dialogue (Du et al., 2023; He et al., 2025a; Ma et al., 2024; Chen et al., 2024; Liang et al., 2023; Jin et al., 2024a; Breum et al., 2024; He et al., 2025c; Salvi et al., 2025). For instance, ChatEval employs a multi-agent referee team to evaluate open-ended responses (Chan et al., 2023), while AgentsCourt improves answer quality through multi-round debate among model instances (He et al., 2024b). Debatrix provides a structured judging framework to assess debates along multiple dimensions (Liang et al., 2024), and DyLAN dynamically assembles agent teams tailored to different tasks (Liu et al., 2024c). Notably, Salvi et al. (2025) shows that GPT-4 equipped with sociodemographic data can outperform humans in persuasion. In this work, we integrate Theory of Mind to strengthen the strategic reasoning and persuasive capacity of LLMs in academic rebuttals, aiming to foster more transparent and constructive scholarly dialogue.
@@
LLM for Academic Peer Review. The emerging paradigm of AI for Research applies Large Language Models (LLMs) to automate and enhance scholarly activities, including automated research (Schmidgall and Moor, 2025; Li et al., 2025; Yamada et al., 2025) and writing assistance (Wang et al., 2025; Chen et al., 2025a). Within the critical domain of peer review, LLMs are leveraged for generating reviews (Zhu et al., 2025; Idahl and Ahmadi, 2025) and for enhancing review quality analysis (Purkayastha et al., 2025). Furthermore, multi-agent systems have been proposed to explore peer review dynamics (Jin et al., 2024b; D’Arcy et al., 2024) and automate research workflows (Schmidgall et al., 2025). Despite the creation of large, multi-turn review datasets (Zhang et al., 2025), there remains limited exploration into the rebuttal stage. Building on these foundations, our work proposes a RebuttalAgent framework that explicitly leverages Theory of Mind to model reviewer intent, enabling more strategic and context-aware responses.
@@ 9 Conclusion
In this paper, we introduce RebuttalAgent, the first framework to ground academic rebuttal in Theory of Mind (ToM), transforming the process into a strategic reasoning task. By integrating evidence-based synthesis with hierarchical reviewer profiling, our model move beyond formulaic templates to facilitate more constructive scholarly dialogue. To support this, we construct RebuttalBench, a large-scale dataset of 70K reasoning chains, and optimize the agent via reinforcement learning with a Self-reward mechanism. Furthermore, our specialized evaluator, Rebuttal-RM, demonstrates superior alignment with human expert preferences compared to proprietary models. Extensive experiments show that RebuttalAgent outperforms base models by 18.3%, achieving performance comparable to advanced models like o3 across both automated and human evaluations.
@@
This framework comprises three main components: (1) a rebuttal evaluator, Rebuttal-RM; (2) a large-scale high-quality dataset, RebuttalBench; and (3) a novel academic assistant, RebuttalAgent. To ensure the full reproducibility of this framework, we have provided detailed documentation across the paper and its appendices. The generation process for the RebuttalBench dataset, along with the complete training procedures for RebuttalAgent (including all hyperparameters), are provided in Section 5. The details for training Rebuttal-RM are provided in Section 6. Our code and models are publicly available at https://github.com/Zhitao-He/RebuttalAgent.
Acknowledgments
This work was supported in part by grants UROP26EG06 and WEB26EG02, Hong Kong University of Science and Technology. We would like to express our sincere gratitude to the anonymous reviewers for their insightful comments and constructive suggestions.